OpenStudy (anonymous):

solve and check. 2x = 15 + x

OpenStudy (anonymous):

You want to get your like terms together. To do that subtract the x to the other side to get: x=15. Then all you have to do is plug that number back into the equation 2x=15+x When you do: 2(15)=15+15 30=30 Thus the x-value works for the equation.

OpenStudy (anonymous):

oh I see--you subtract x from both sides of the equation?